Retries 1 If n=11, ¯ x (x-bar)=46, and s=2, find the margin of error at a 99% confidence level (use at least three decimal places)
Answer #1

We have given that,              
              
Sample mean =46      
Sample standard deviation =2  
Sample size =11      
Level of significance=1-0.99=0.01      
Degree of freedom =10      
              
t critical value is (by using t table)=   3.169      
E= t*\frac{S}{\sqrt{n}}= 3.169 *(\frac { 2 } {\sqrt{ 11 }})=1.912

Margin of error =1.912